If x+(1/y)=1 and y+(1/z)=1, what is the value of the product xyz?
Answer by scott8148(6628)

x+(1/y)=1 ___ xy + 1 = y

substituting ___ xy + 1 + 1/z = 1 ___ xyz + z + 1 = z ___ xyz = -1